# Critical exponents of the Ising model with quenched structural disorder and long-range interactions at spatial dimension $d=3$

### v2

@@ 3 RG approach
The expression (3.7) is exact since $Z_{\phi}=1$, because LR term ($k^{\sigma}$) does not renormalize via its nonanalyticity in momentum spase [13, 14]. Therefore pair correlation critical exponent is $\eta=2-\sigma$ for the cases where LR interactions are relevant for critical behaviour. Such result for $\eta$ is supported by other approaches (see review in [18]). Here $\epsilon=2\sigma-d$ and notations in (3.4)-(3.6) are following: $D_{\sigma}=\psi(1)-2\psi\left(\frac{\sigma}{2}\right)+\psi(\sigma)$, $E_{\sigma}=\psi^{(1)}(\sigma)-\psi^{(1)}\left(\frac{\sigma}{2}\right)$}, $F_{\sigma}=\psi^{(1)}(1)-\psi^{(1)}(\sigma)$}, $G_{\sigma}=\frac{\Gamma\left(-\frac{\sigma}{2}\right)\Gamma(\sigma)^{2}}{\Gamma\left(\frac{3\sigma}{2}\right)}$; where $\Gamma(\dots)$ is the gamma function, while $\psi_{i}(\dots)$ is the digamma function of $i$-order. Results for $H_{\sigma}$ was obtained only numerically (for details in computation and numerical estimation for certain $\sigma$, see Appendix A). At the two-loop order (first lines of Eqs. (3.4)-(3.6)) we reproduce known RG results up to normalization constants [30].
Figure 2: Schematic RG flows (denoted by red arrows) for the LR interacting random Ising model at $d<2\sigma$ in the parametric space of coupling constants $(u,v)$. Unstable FPs $G$ and $P$ are shown by blue discs, while stable FPs UUand $R$ are depicted yellow squares.
Similarly to the SR case [42], the $\beta$-functions (3.4), (3.5) for LR models with quenched dilution are also known to have four FPs $(u^{*},\,v^{*})$ in the parameter space $(u,v)$ [28, 32, 33]. The physically relevant FPs are those that correspond to the initial conditions of the effective Hamiltonian (2.8), namely $u>0$ and $v<0$. For our case in dimension $d<2\sigma$ we have the following picture (see Fig. 2): (i) unstable Gaussian FP G{G} ($u^{*}=0$, $v^{*}=0$); (ii) stable unphysical FP U{U} ($u^{*}=0$, $v^{*}\neq 0$), but it is inaccessible from the physical initial conditions $u>0$ and $v<0$ set by (2.8); (iii) unstable pure (system without impurities) FP P{P} ($u^{*}\neq 0$, $v^{*}=0$); (iv) stable random (disorder-induced) FP R{R} ($u^{*}\neq 0$, $v^{*}\neq 0$). Therefore only FP $R$ corresponds to the critical point of the LR interacting Ising model with quenched dilution. The calculation of the critical exponent $\nu$ for this fixed point is presented in the next Section.
@@ 4.1 Resummation procedure
$f(z)=\int_{0}^{\infty}e^{-t}\,t^{b}\,B(zt)\,dt.$ (4.2)
In the Padé-Borel-Leroy resummation procedure Borel image (4.1) in (4.2) is replaced by its Padé approximant. In our calculation we use Padé-approximant [L1/1](zt)[L-1/1](zt) with linear denominator:
$f^{res}(z)=\int_{0}^{\infty}e^{-t}\,t^{b}\,[L-1/1](zt)\,dt.$ (4.3)
For $b=0$, (4.3) reduces to the standard Padé–Borel resummation [56, 57]. Let us note that any converged observable should be independent of artificially introduced parameters, such as the Leroy parameter. However, at any finite loop order $L$, the resummed series acquire an artificial dependence on such parameters. To extract a meaningful result, we assume that the optimal estimate of an observable at order $L$ corresponds to the values of the artificial parameters for which the resummed quantity varies most weakly – that is, where it becomes stationary with respect to changes in these parameters. This criterion is known as the Principle of Minimal Sensitivity [58, 59]. Nevertheless, even the Padé–Borel–Leroy method may fail for multivariable series due to poles lying on the integration contour or poor convergence of Padé approximants.
Generalisation of the Padé-Borel-Leroy procedure on the case of several variables is obvious with help of a resolvent series in an auxiliary variable. For instance, writing any function $f(u,v)$ as $f(uz,vz)$, one can perform procedure (4.1)-(4.3) for the series in $z$, eventually setting $z=1$ after all steps.
@@ 4.2 Results at fixed dimension $d=3$
Let us start the discussion of the values of the correlation length critical exponent $\nu$ for the physically relevant case $d=3$, considering the two-loop RG functions (first lines of Eqs. (3.4)-(3.6)). We provide the intermediate technical details in the Appendix B. Applying the Padé-Borel-Leroy resummation to the two-loop $\beta_{u}$ and $\beta_{v}$ functions, we are able to found the FP $R$ by fixing $d=3$ for all values $\sigma$ between $3/2$ and $2-\eta_{SR}$, where we use the result $\eta_{SR}=0.0362$ calculated for the SR Ising model in the highest accessible loop order [60]. At $\sigma=3/2$ (case of upper critical dimension for three dimensional case, $\sigma=d/2$) the logarithmic corrections to the scaling are expected in the same way as in SR theory at $d=4$, however we do not touch this issue in our study.
@@
Substituting the obtained FP coordinates into the correspondingly resummed expression for the critical exponent $\nu$ we obtain data shown by red solid line in Fig. 3. The results are obtained for the optimal Leroy parameter, $b_{\mathrm{opt}}=27$ (see Appendix B). The dependence of $\nu$ on $\sigma$ is monotonic with a local maximum at $\sigma\approx 1.55$. With the increase of $\sigma$, the values of $\nu$ go towards the value of the correlation length critical exponent of the SR random Ising model, $\nu_{RSR}$. We present its value $\nu_{RSR}=0.665$, obtained applying the same Pad-́Borel-Leroy resummation to the corresponding SR RG functions, restricted at two-loop level, by the dashed red line in Fig. 3. The value $\nu_{RSR}=0.675$ calculated from RG functions accessible in the highest loop-approximation [41] is shown by the orange dot-dashed line.
@@
For comparison, we present in Fig. 3 the two-loop results for pure (undiluted) LR Ising model by dot-dashed line. They are obtained by extracting data for FP $P$ by solving the equation $\beta_{u}(u^{*},0)=0$ with the correspondingly resummed function $\beta_{u}$ and substituting into the expression for $\nu$. The significant difference between the critical exponents of the random LR Ising model and those of its undiluted counterpart is observed. Therefore, it is interesting to verify this finding using another approach.
@@
The previously used simple resummation techniques, which gave results for two-loop RG functions, turn out to be insufficiently reliable when applied to the fixed points derived from the three-loop RG functions (see discussion in the Appendix B). However, we cannot use more sophisticated methods based on conformal mapping [58, 59], since series for our RG functions obtained in low-loop approximations are two short.
@@
Figure 3: Correlation length exponent $\nu$ for the Ising model with LR interactions and quenched disorder at $d=3$ as a function of $\sigma$. The data in the region $d/2<\sigma<2-\eta_{SR}$ ( denoted as RLR) are obtained within the two-loop approximation with the help of the Padé–Borel–Leroy procedure at $b_{\mathrm{opt}}=27$ (red solid curve) and within the three-loop approximation with the help of the “truncated” version of the subsequent resummation procedure (solid black curve). The values of $\nu$ for the mean field (MF) region ($\sigma\leq d/2$) are shown by thin dashed line. Estimates of $\nu$ for the random SR (RSR) model, valid for $\sigma<2-\eta_{SR}$, obtained within different approximations, are also shown: two-loop approximation value $\nu_{RSR}=0.665$ (red dashed line), three-loop approximation value $\nu_{RSR}=0.668$ (black dashed line), six loop approximation value $\nu_{RSR}=0.675$ [41]. The light red dot-dashed curve presents two-loop data for the critical exponent $\nu$ of the undiluted LR Ising model.
@@
Therefore, we employ a subsequent resummation procedure which has been shown to perform well in studies of the three-dimensional SR Ising model with weak quenched disorder. This method is designed for the resummation of a two-variable asymptotic series arising in the RG analysis of systems with quenched disorder [61, 62, 63]. The technique involves the consecutive application of the Padé–Borel–Leroy transformation: first with respect to the coupling associated with the pure system ($u$), and then with respect to the disorder-induced coupling ($v$), thereby implementing the principle of “inequality of rights” [64] between variables (for more details, see Appendix C). This asymmetrical resummation procedure may help to avoid Padé poles and provides stable estimates for the fixed points and their stability within high-order perturbative expansions [62, 63]. The most reliable fixed point estimates extracted from the three-loop RG functions are provided by the “truncated” version of the subsequent resummation procedure, which remains more stable under variations of the control parameter $\sigma$ (black solid line in Fig. 3).
@@
As one can see, for small values of $\sigma<1.75$ the three-loop results are very close to the two-loop values. With a further increase of $\sigma$, a larger deviation of the $\nu$ values from different approximations is observed. After $\sigma>1.9$ the three loop data for $\nu$ abruptly go down, which does not allow to obtain a reliable estaimate. Such behaviour raises the question of whether the two-loop approximation should be considered the optimal truncation for the random LR Ising model, or even of the validity of the perturbative RG approach in this case.
@@ 5 Conclusion
In this work, we investigated how the critical behaviour of an Ising system changes under the mutual influence of two factors: LR interactions and structural disorder. It is known that the random Ising model with LR belongs to a new “random” universality class, distinct from the pure (undiluted) LR universality class, as well as from the random SR universality class for the values of decay parameter $d/2<\sigma<2-\eta_{SR}$.
Aiming to provide a quantitative description of the critical behaviour in this new universality class, we considered a scalar $\phi^{4}$ field theory with two coupling constants and employed the field-theoretic RG approach. We have derived the three-loop RG functions within massless minimal subtraction scheme. On the basis of these functions, we have calculated the correlation length critical exponent $\nu$ at fixed spatial dimension $d=3$ and for the values of decay parameter $d/2<\sigma<2-\eta_{SR}$ employing various techniques for analysing asymptotic series.
Our two-loop estimates obtained with help Padé–Borel–Leroy resummation methods demonstrate a non-monotonic dependence of $\nu$ on $\sigma$ when changing from mean-field boundary $d/2$ to the SR boundary $2-\eta_{SR}$. There is a significant difference with the corresponding values of $\nu$ for the undiluted LR Ising model. Therefore, we hope that our study will motivate verification of this finding within Monte Carlo simulations.
While within the three-loop approximation we have obtained reliable estimates for $\nu$ using the subsequent resummation approach only for $\sigma<1.9$. A possible explanation of this outcome is that the perturbative field-theoretical RG approach is not proper for the random Ising model with LR interactions, making the two-loop approximation to be optimal truncation in this case.
@@ Acknowledgements
We thank Prof. Yu. Holovatch for the introduction us to the fascinating world of critical phenomena and phase transitions, and we acknowledge numerous useful discussions with him during the work on this topic. The work was supported by the National Research Foundation of Ukraine Project 2023.03/0099 “Criticality of complex systems: fundamental aspects and applications”.
@@ Appendix A Computation $H_{\sigma}$
The contribution $H_{\sigma}$ in (3.4), (3.5) comes from the three-loop integral depicted by diagram $I_{4}$ in [19, 24] (see inset of Fig. 4). As it was shown in [24] it reduces to calculation of kite diagram. Here, we employ the same method [24], which is based on a decomposition into Hepp sectors, i.e. an ordering of Schwinger parameters that isolates $1/\epsilon$ divergences. Considering $I_{4}$ in the Schwinger-parameter representation, we obtain:
$I_{4}\sim\int_{0}^{\infty}d\bar{a}_{1}\ldots d\bar{c}_{2}\frac{(\bar{a}_{1}\bar{a}_{2}\bar{b}_{1}\bar{b}_{2}\bar{c}_{1}\bar{c}_{2})^{\sigma/2-1}e^{-(\bar{a}_{1}+\bar{a}_{2}+\bar{b}_{1}+\bar{b}_{2}+\bar{c}_{1}+\bar{c}_{2})}}{S^{d/2}}$ (A.1)
@@
Here, as mention in [24], the denominator $S$ is symmetric by (i) exchange of the pairs $(\bar{a}_{1},\bar{a}_{2})$$(\bar{b}_{1},\bar{b}_{2})$, $(\bar{c}_{1},\bar{c}_{2})$ and by (ii) simultaneous exchange of elements of two of the pairs. As a result, instead of the full $6!=720$ sectors we have $(6!)/(3!2^{2})=30$ distinct sectors need to be considered.
@@
$H_{\sigma}=48\frac{\Gamma{\left(\sigma\right)^{3}}}{\Gamma{\left(\frac{\sigma}{2}\right)^{6}}}\sum_{j=1}^{30}\int_{0}^{1}\prod_{i=1}^{5}dt_{i}\frac{(t_{1}t_{5})^{\sigma/2-1}(t_{2}t_{4})^{\sigma-1}(t_{3})^{3\sigma/2-1}}{\tilde{S}_{j}^{\sigma}},$ (A.3)
which can be evaluated numerically for specific values of $\sigma$ (see Fig. 4). Each $\tilde{S}_{j}$ in (A.3) is the sector-specific polynomial that arises after (i) ordering the six Schwinger parameters into one of the 30 inequivalent sectors, (ii) performing the change of variables $\{\bar{a}_{1},\bar{a}_{2},\bar{b}_{1},\bar{b}_{2},\bar{c}_{1},\bar{c}_{2}\}\to\{\bar{a}_{1},t_{1},t_{2},t_{3},t_{4},t_{5}\}$ and integrating out $\bar{a}_{1}$.
Figure 4: Numerical evaluation of $H_{\sigma}$ in eq. (A.3) with respect to the control parameter $\sigma$. The numerical integration error is estimated as the difference between results obtained at different levels of computational refinement and is determined by the extent to which the outcome stabilises upon adding additional sampling points. The resulting uncertainty remains within the size of the plotted markers and does not affect the qualitative behaviour of the curves.

### v3

@@ 3 RG approach
The expression (3.7) is exact since $Z_{\phi}=1$, because LR term ($k^{\sigma}$) does not renormalize via its nonanalyticity in momentum spase [13, 14]. Therefore, pair correlation critical exponent is $\eta=2-\sigma$ for the cases where LR interactions are relevant for critical behaviour. Such a result for $\eta$ is supported by other approaches (see review in [18]). Here, $\epsilon=2\sigma-d$ and notations in (3.4)–(3.6) are as follows: $D_{\sigma}=\psi(1)-2\psi\left(\frac{\sigma}{2}\right)+\psi(\sigma)$, $E_{\sigma}=\psi^{(1)}(\sigma)-\psi^{(1)}\left(\frac{\sigma}{2}\right)$}, $F_{\sigma}=\psi^{(1)}(1)-\psi^{(1)}(\sigma)$}, $G_{\sigma}=\frac{\Gamma\left(-{\sigma}/{2}\right)\Gamma(\sigma)^{2}}{\Gamma\left({3\sigma}/{2}\right)}$; where $\Gamma(\dots)$ is the gamma function, while $\psi^{(i)}(\dots)$ is the digamma function of $i$-order. Results for $H_{\sigma}$ were obtained only numerically (for details in computation and numerical estimation for certain $\sigma$, see appendix A). At the two-loop order (first lines of equations (3.4)–(3.6)) we reproduce the known RG results up to normalization constants [30].
Figure 2: (Colour online) Schematic RG flows (denoted by red arrows) for the LR interacting random Ising model at $d<2\sigma$ in the parametric space of coupling constants $(u,v)$. Unstable FPs $G$ and $P$ are shown by blue discs, while stable FPs UUand $R$ are depicted with yellow squares.
Similarly to the SR case [42], the $\beta$-functions (3.4), (3.5) for LR models with quenched dilution are also known to have four FPs $(u^{*},\,v^{*})$ in the parameter space $(u,v)$ [28, 32, 33]. The physically relevant FPs are those that correspond to the initial conditions of the effective Hamiltonian (2.8), namely $u>0$ and $v<0$. For our case in dimension $d<2\sigma$, we have the following picture (see figure 2): (i) unstable Gaussian FP G{G} ($u^{*}=0$, $v^{*}=0$); (ii) stable unphysical FP U{U} ($u^{*}=0$, $v^{*}\neq 0$), but it is inaccessible from the physical initial conditions $u>0$ and $v<0$ set by (2.8); (iii) unstable pure (system without impurities) FP P{P} ($u^{*}\neq 0$, $v^{*}=0$); (iv) stable random (disorder-induced) FP R{R} ($u^{*}\neq 0$, $v^{*}\neq 0$). Therefore, only FP $R$ corresponds to the critical point of the LR interacting Ising model with quenched dilution. The calculation of the critical exponent $\nu$ for this fixed point is presented in the next section.
@@ 4.1 Resummation procedure
$f(z)=\int_{0}^{\infty}\mathrm{e}^{-t}\,t^{b}\,B(zt)\,\mathrm{d}t.$ (4.2)
In the Padé–Borel–Leroy resummation procedure Borel image (4.1) in (4.2) is replaced by its Padé approximant. In our calculation we use Padé-approximant [L1/1](zt)[L-1/1](zt) with linear denominator:
$f^{\rm res}(z)=\int_{0}^{\infty}\mathrm{e}^{-t}\,t^{b}\,[L-1/1](zt)\,\mathrm{d}t.$ (4.3)
For $b=0$, (4.3) reduces to the standard Padé–Borel resummation [56, 57]. Let us note that any converged observable should be independent of artificially introduced parameters, such as the Leroy parameter. However, at any finite loop order $L$, the resummed series acquire an artificial dependence on such parameters. To extract a meaningful result, we assume that the optimal estimate of an observable at order $L$ corresponds to the values of the artificial parameters for which the resummed quantity varies most weakly — that is, where it becomes stationary with respect to the changes in these parameters. This criterion is known as the Principle of Minimal Sensitivity [58, 59]. Nevertheless, even the Padé–Borel–Leroy method may fail for multivariable series due to the poles lying on the integration contour or due to poor convergence of Padé approximants.
Generalization of the Padé–Borel–Leroy procedure on the case of several variables is obvious with the help of a resolvent series in an auxiliary variable. For instance, writing any function $f(u,v)$ as $f(uz,vz)$, one can perform procedure (4.1)–(4.3) for the series in $z$, eventually setting $z=1$ after all steps.
@@ 4.2 Results at fixed dimension $d=3$
Let us start the discussion of the values of the correlation length critical exponent $\nu$ for the physically relevant case $d=3$, considering the two-loop RG functions (first lines of equations (3.4)–(3.6)). We provide the intermediate technical details in the appendix B. Applying the Padé–Borel–Leroy resummation to the two-loop $\beta_{u}$ and $\beta_{v}$ functions, we are able to found the FP $R$ by fixing $d=3$ for all values $\sigma$ between $3/2$ and $2-\eta_{\rm SR}$, where we use the result $\eta_{\rm SR}=0.0362$ calculated for the SR Ising model in the highest accessible loop order [60]. At $\sigma=3/2$ (the case of upper critical dimension for three dimensional case, $\sigma=d/2$) the logarithmic corrections to the scaling are expected in the same way as in SR theory at $d=4$, although we do not touch this issue in our study.
@@
Substituting the obtained FP coordinates into the correspondingly resummed expression for the critical exponent $\nu$, we obtain the data shown with red solid line in figure 3. The results are obtained for the optimal Leroy parameter, $b_{\mathrm{opt}}=27$ (see appendix B). The dependence of $\nu$ on $\sigma$ is monotonic with a local maximum at $\sigma\approx 1.55$. With an increase of $\sigma$, the values of $\nu$ go towards the value of the correlation length critical exponent of the SR random Ising model, $\nu_{\rm RSR}$. We present its value $\nu_{\rm RSR}=0.665$, obtained applying the same Padé–Borel–Leroy resummation to the corresponding SR RG functions, restricted at two-loop level, by the dashed red line in figure 3. The value $\nu_{\rm RSR}=0.675$ calculated from RG functions accessible in the highest loop-approximation [41] is shown with the orange dot-dashed line.
@@
For comparison, we present in figure 3 the two-loop results for pure (undiluted) LR Ising model by dot-dashed line. They are obtained by extracting data for FP $P$ by solving the equation $\beta_{u}(u^{*},0)=0$ with the correspondingly resummed function $\beta_{u}$ and substituting into the expression for $\nu$. The significant difference between the critical exponents of the random LR Ising model and those of its undiluted counterpart is observed. Therefore, it is interesting to verify this finding using another approach.
@@
The previously used simple resummation techniques, which gave results for two-loop RG functions, turn out to be insufficiently reliable when applied to the fixed points derived from the three-loop RG functions (see discussion in the appendix B). However, we cannot use more sophisticated methods based on conformal mapping [58, 59], since the series for our RG functions obtained in low-loop approximations are two short.
@@
Figure 3: (Colour online) Correlation length exponent $\nu$ for the Ising model with LR interactions and quenched disorder at $d=3$ as a function of $\sigma$. The data in the region $d/2<\sigma<2-\eta_{\rm SR}$ (denoted as RLR) are obtained within the two-loop approximation with the help of the Padé–Borel–Leroy procedure at $b_{\mathrm{opt}}=27$ (red solid curve) and within the three-loop approximation with the help of the “truncated” version of the subsequent resummation procedure (solid black curve). The values of $\nu$ for the mean field (MF) region ($\sigma{\leqslant}d/2$) are shown by thin dashed line. Estimates of $\nu$ for the random SR (RSR) model, valid for $\sigma<2-\eta_{\rm SR}$, obtained within different approximations, are also shown: two-loop approximation value $\nu_{\rm RSR}=0.665$ (red dashed line), three-loop approximation value $\nu_{\rm RSR}=0.668$ (black dashed line), six loop approximation value $\nu_{\rm RSR}=0.675$ [41]. The light red dot-dashed curve presents two-loop data for the critical exponent $\nu$ of the undiluted LR Ising model.
@@
Therefore, we employ a subsequent resummation procedure which has been shown to perform well in studies of the three-dimensional SR Ising model with weak quenched disorder. This method is designed for the resummation of a two-variable asymptotic series arising in the RG analysis of systems with quenched disorder [61, 62, 63]. The technique involves the consecutive application of the Padé–Borel–Leroy transformation: first with respect to the coupling associated with the pure system ($u$), and then with respect to the disorder-induced coupling ($v$), thereby implementing the principle of “inequality of rights” [64] between variables (for more details, see appendix C). This asymmetrical resummation procedure may help to avoid Padé poles and provides stable estimates for the fixed points and their stability within high-order perturbative expansions [62, 63]. The most reliable fixed point estimates extracted from the three-loop RG functions are provided by the “truncated” version of the subsequent resummation procedure, which remains more stable under variations of the control parameter $\sigma$ (black solid line in figure 3).
@@
As one can see, for small values of $\sigma<1.75$, the three-loop results are very close to the two-loop values. With a further increase of $\sigma$, a larger deviation of the $\nu$ values from different approximations is observed. After $\sigma>1.9$ the three loop data for $\nu$ abruptly go down, which does not allow to obtain a reliable estaimate. Such behaviour raises the question of whether the two-loop approximation should be considered the optimal truncation for the random LR Ising model, or even of the validity of the perturbative RG approach in this case.
@@ 5 Conclusion
In this work, we investigated how the critical behaviour of an Ising system changes under the mutual influence of two factors: LR interactions and structural disorder. It is known that the random Ising model with LR belongs to a new “random” universality class, distinct from the pure (undiluted) LR universality class, as well as from the random SR universality class for the values of decay parameter $d/2<\sigma<2-\eta_{\rm SR}$.
Aiming at providing a quantitative description of the critical behaviour in this new universality class, we considered a scalar $\phi^{4}$ field theory with two coupling constants and employed the field-theoretic RG approach. We have derived the three-loop RG functions within massless minimal subtraction scheme. On the basis of these functions, we have calculated the correlation length critical exponent $\nu$ at a fixed spatial dimension $d=3$ and for the values of decay parameter $d/2<\sigma<2-\eta_{\rm SR}$ employing various techniques for analysing asymptotic series.
Our two-loop estimates obtained with help Padé–Borel–Leroy resummation methods demonstrate a non-monotonous dependence of $\nu$ on $\sigma$ when changing from the mean-field boundary $d/2$ to the SR boundary $2-\eta_{\rm SR}$. There is a significant difference with the corresponding values of $\nu$ for the undiluted LR Ising model. Therefore, we hope that our study will motivate verification of this finding within Monte Carlo simulations.
Herein, within the three-loop approximation we have obtained reliable estimates for $\nu$ using the subsequent resummation approach only for $\sigma<1.9$. A possible explanation of this outcome is that the perturbative field-theoretical RG approach is not proper for the random Ising model with LR interactions, making the two-loop approximation to be optimal truncation in this case.
@@ Acknowledgements
We thank Prof. Yu. Holovatch for having introduced us to the fascinating world of critical phenomena and phase transitions, and we acknowledge numerous useful discussions with him during the work on this topic. The work was supported by the National Research Foundation of Ukraine Project 2023.03/0099 “Criticality of complex systems: fundamental aspects and applications”.
@@ Appendix A Computation $H_{\sigma}$
The contribution $H_{\sigma}$ in (3.4), (3.5) comes from the three-loop integral depicted by diagram $I_{4}$ in [19, 24] (see inset of figure A.1). As it was shown in [24] it reduces to calculation of kite diagram. Here, we employ the same method [24], which is based on a decomposition into Hepp sectors, i.e., an ordering of Schwinger parameters that isolates $1/\epsilon$ divergences. Considering $I_{4}$ in the Schwinger-parameter representation, we obtain:
$I_{4}\sim\int_{0}^{\infty}\mathrm{d}\bar{a}_{1}\ldots\mathrm{d}\bar{c}_{2}\frac{(\bar{a}_{1}\bar{a}_{2}\bar{b}_{1}\bar{b}_{2}\bar{c}_{1}\bar{c}_{2})^{\sigma/2-1}\mathrm{e}^{-(\bar{a}_{1}+\bar{a}_{2}+\bar{b}_{1}+\bar{b}_{2}+\bar{c}_{1}+\bar{c}_{2})}}{S^{d/2}}$ (A.1)
@@
Here, as mentioned in [24], the denominator $S$ is symmetric by (i) exchange of the pairs $(\bar{a}_{1},\bar{a}_{2})$$(\bar{b}_{1},\bar{b}_{2})$, $(\bar{c}_{1},\bar{c}_{2})$ and by (ii) simultaneous exchange of elements of two of the pairs. As a result, instead of the full $6!=720$ sectors we have $(6!)/(3!2^{2})=30$ distinct sectors need to be considered.
Figure A.1: (Colour online) Numerical evaluation of $H_{\sigma}$ in equation (A.3) with respect to the control parameter $\sigma$. The numerical integration error is estimated as the difference between the results obtained at different levels of computational refinement and is determined by the extent to which the outcome stabilizes upon adding additional sampling points. The resulting uncertainty remains within the size of the plotted markers and does not affect the qualitative behaviour of the curves.
@@
$H_{\sigma}=48\frac{\Gamma{\left(\sigma\right)^{3}}}{\Gamma{\left(\frac{\sigma}{2}\right)^{6}}}\sum_{j=1}^{30}\int_{0}^{1}\prod_{i=1}^{5}\mathrm{d}t_{i}\frac{(t_{1}t_{5})^{\sigma/2-1}(t_{2}t_{4})^{\sigma-1}(t_{3})^{3\sigma/2-1}}{\tilde{S}_{j}^{\sigma}},$ (A.3)
which can be evaluated numerically for specific values of $\sigma$ (see figure A.1). Each $\tilde{S}_{j}$ in (A.3) is the sector-specific polynomial that arises after (i) ordering the six Schwinger parameters into one of the 30 inequivalent sectors, (ii) performing the change of variables $\{\bar{a}_{1},\bar{a}_{2},\bar{b}_{1},\bar{b}_{2},\bar{c}_{1},\bar{c}_{2}\}\to\{\bar{a}_{1},t_{1},t_{2},t_{3},t_{4},t_{5}\}$ and integrating out $\bar{a}_{1}$.